Solve for the unknown interest rate in each of the following: (Do not round intermediate calculations and enter your answersSolve for the unknown number of years in each of the following: (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

a). r = [FV / PV]1/n - 1

= [$1,441 / $745]1/4 - 1

= [1.9342]1/4 - 1

= 1.1793 - 1 = 0.1793, or 17.93%

b). r = [FV / PV]1/n - 1

= [$1,778 / $935]1/5 - 1

= [1.9016]1/5 - 1

= 1.1372 - 1 = 0.1372, or 13.72%

c). r = [FV / PV]1/n - 1

= [$144,832 / $18,000]1/16 - 1

= [8.0462]1/16 - 1

= 1.1392 - 1 = 0.1392, or 13.92%

d). r = [FV / PV]1/n - 1

= [$318,815 / $73,300]1/19 - 1

= [4.3495]1/19 - 1

= 1.0804 - 1 = 0.0804, or 8.04%